# BRiTE: Bootstrapping Reinforced Thinking Process to Enhance Language Model Reasoning

### v1

@@ 3.2 Bootstrapping Reinforced Thinking Process
This lemma is equivalent to the non-negativity of the KL divergence. For a detailed proof, please refer to Appendix A.1. ∎
@@
Suppose Assumption 3.2 holds. Suppose Assumption 3.2 holds. Given that $\mathcal{L}$ is concave and $\theta^{*}=\mathop{\mathrm{argmax}}_{\theta}\mathcal{L}(\theta)$, we have
@@
See Appendix B.1 for a detailed proof. ∎
Theorem 3.3 establishes a convergence rate of $1/T$, where $T$ represents the number of iteration steps. In our analysis, we link our algorithm to classical mirror descent algorithms (Nemirovskij and Yudin, 1983; Bubeck et al., 2015), which have also been utilized in recent studies of policy optimization algorithms (Agarwal et al., 2021; Cai et al., 2020; Zhong and Zhang, 2024). Additionally, our results rely on a concave assumption. This assumption is crucial for proving global convergence results, and we can still obtain a weaker guarantee for stationary points even without this concave condition; see Appendix B.2 for details.
@@ 3.4 Practical Implementation: The Power of Reinforcement Learning
See Appendix A.2 for a detailed proof. ∎
@@ 4.1 Experimental Setups
    BRiTE-SFT first obtains $\mathbb{Q}$ in (3.4) through RL, then updates $\theta$ in (3.2) using SFT. The SFT data consists of three components: problems $x$, thinking processes $z_{\mathbb{Q}}(x)$ generated by $\mathbb{Q}$, along with the ground truth answers $y^{*}$.
@@ 4.2 Experimental Result and Analysis
Finally, we extend the evaluation of BRiTE beyond mathematics tasks to assess its performance on code generation. The results, presented in Table 2, demonstrate a consistent trend observed in mathematics tasks: BRiTE outperforms rejection sampling-based algorithms and even surpasses SFT using human-annotated answers. This highlights BRiTE’s versatility and effectiveness across diverse problem domains, showcasing its ability to generalize to a wide range of reasoning tasks. We also remark that rejection sampling-based algorithms require the code datasets to be equipped with correct unit tests, which are unnecessary for BRiTE. For a detailed description of the experimental setup and configuration, refer to Appendix C.2.
@@ Acknowledgments
Appendix A Missing Proofs in the Main Paper
A.1 Proof of Lemma 3.1
@@ Appendix A Missing Proofs in the Main Paper
$\mathbb{E}_{w\sim q(\cdot)}[\log P_{w}-\log q(w)]-\log\sum_{w\in\mathcal{W}}P_{z}$ $=\mathbb{E}_{w\sim q(\cdot)}\Big{[}\log\frac{P_{w}}{\sum_{w^{\prime}\in\mathcal{W}}P_{w^{\prime}}}-\log q(w)\Big{]}=-\mathrm{KL}(q\|\widetilde{p})\leq 0,$
@@
Appendix B Convergence Results
B.1 Proof of Theorem 3.3
@@ Appendix B Convergence Results
$=A(x,\theta^{\prime})-A(x,\theta)+\left\langle\mathbb{E}_{(z,y)\sim\mathbb{P}(\cdot,\cdot{\,|\,}x,\theta)}[\mathcal{K}((x,z,y),\cdot)],f_{\theta}-f_{\theta^{\prime}}\right\rangle,$ (B.1)
@@
$=\sum_{(z,y)\in\mathcal{Z}\times\mathcal{Y}}\mathbb{P}(z,y{\,|\,}x,\theta)\cdot\mathcal{K}((x,z,y),\cdot),$ (B.2)
where $\nabla$ is a functional gradient. Combining (B.1) and (B.1), we have
@@
which finishes the proof of Lemma B.1. ∎
Lemma B.2.
@@
$=\frac{\sum_{(z,y,o)\in\mathscr{Z}\times\mathscr{Y}\times\mathscr{O}}\big{(}\mathbb{P}(o{\,|\,}x,z,y)\cdot\nabla\mathbb{P}(z,y{\,|\,}x,\theta)\big{)}}{\sum_{(z,y,o)\in\mathscr{Z}\times\mathscr{Y}\times\mathscr{O}}\mathbb{P}(z,y,o{\,|\,}x,\theta)},$ (B.3)
@@
$=\mathbb{P}(z,y{\,|\,}x,\theta)\cdot[\mathcal{K}((x,z,y),\cdot)-\nabla A(x,\theta)].$ (B.4)
Combining (B.1) and (B.1), we have
@@
$=\mathbb{E}_{(z,y)\sim\mathbb{Q}(\cdot,\cdot{\,|\,}x,\theta)}[\mathcal{K}((x,z,y),\cdot)]-\nabla A(x,\theta).$ (B.5)
This finishes the proof of Lemma B.2. ∎
@@
$=\mathop{\mathrm{argmax}}_{\theta}\Big{\{}\left\langle\mathbb{E}_{(z,y)\sim\mathbb{Q}(\cdot,\cdot{\,|\,}x,\theta_{t})}[\mathcal{K}((x,z,y),\cdot)],f_{\theta}-f_{\theta_{t}}\right\rangle+A(x,\theta_{t})-A(x,\theta)\Big{\}},$ (B.6)
@@
Combining (B.1) and Lemma B.2, we have
$\theta_{t+1}$ $=\mathop{\mathrm{argmax}}_{\theta}\Big{\{}\left\langle\nabla\log\mathbb{P}(z\in\mathscr{Z},y\in\mathscr{Y},o\in\mathscr{O}{\,|\,}x,\theta_{t})+\nabla A(x,\theta_{t}),f_{\theta}-f_{\theta_{t}}\right\rangle+A(x,\theta_{t})-A(x,\theta)\Big{\}}.$ (B.7)
@@
$\nabla A(\theta_{t+1})-\nabla A(\theta_{t})=\nabla\log\mathbb{P}(z\in\mathscr{Z},y\in\mathscr{Y},o\in\mathscr{O}{\,|\,}x,\theta_{t}).$ (B.8)
@@
where the last equality follows from Lemma B.1. This is equivalent to
@@
$\qquad=\underbrace{\langle\nabla\log\mathbb{P}(z\in\mathscr{Z},y\in\mathscr{Y},o\in\mathscr{O}{\,|\,}x,\theta_{t}),f_{\theta^{*}}-f_{\theta_{t+1}}\rangle}_{(\star)}+\mathrm{KL}\big{(}\mathbb{P}(\cdot,\cdot{\,|\,}x,\theta_{t})\|\mathbb{P}(\cdot,\cdot{\,|\,}x,\theta_{t+1})\big{)},$ (B.9)
where the last equality follows from (B.8). For the Term ()($\star$)( ⋆ ), we have
@@
$=\underbrace{\left\langle\nabla\log\mathbb{P}(z\in\mathscr{Z},y\in\mathscr{Y},o\in\mathscr{O}{\,|\,}x,\theta_{t}),f_{\theta^{*}}-f_{\theta_{t}}\right\rangle}_{\mathrm{(I)}}+\underbrace{\left\langle\nabla\log\mathbb{P}(z\in\mathscr{Z},y\in\mathscr{Y},o\in\mathscr{O}{\,|\,}x,\theta_{t}),f_{\theta_{t}}-f_{\theta_{t+1}}\right\rangle}_{\mathrm{(II)}}.$ (B.10)
@@
$\log\mathbb{P}(z\in\mathscr{Z},y\in\mathscr{Y},o\in\mathscr{O}{\,|\,}x,\theta_{t+1})$ (B.12)
@@
$\mathbb{Q}(z,y,o{\,|\,}x,\theta)=\frac{\mathbb{P}(z,y,o{\,|\,}x,\theta)}{\sum_{(z^{\prime},y^{\prime},o^{\prime})\in\mathscr{Z}\times\mathscr{Y}\times\mathscr{O}}\mathbb{P}(z^{\prime},y^{\prime},o^{\prime}{\,|\,}x,\theta)}=\frac{\mathbb{P}(z,y,o{\,|\,}x,\theta)}{\mathbb{P}(z\in\mathscr{Z},y\in\mathscr{Y},o\in\mathscr{O}{\,|\,}x,\theta)}.$ (B.13)
@@
$=\sum_{(z,y,o)\in\mathscr{Z}\times\mathscr{Y}\times\mathscr{O}}\mathbb{Q}(z,y,o{\,|\,}x,\theta_{t})\cdot\log\frac{\mathbb{P}(z,y,o{\,|\,}x,\theta_{t})}{\mathbb{Q}(z,y,o{\,|\,}x,\theta_{t})},$ (B.14)
where the last equality is implied by the definition of $\mathbb{Q}(z,y,o{\,|\,}x,\theta)$ in (B.13). Combining (LABEL:eq:l:t+1) and (B.1), we have
@@
$\qquad=\sum_{(z,y,o)\in\mathscr{Z}\times\mathscr{Y}\times\mathscr{O}}\mathbb{Q}(z,y,o{\,|\,}x,\theta_{t})\cdot\log\frac{\mathbb{P}(z,y{\,|\,}x,\theta_{t+1})}{\mathbb{P}(z,y{\,|\,}x,\theta_{t})},$ (B.15)
where the third equality follows from the non-negativity of KL-divergence, and the last equality is implied by (3.1). By Assumption 3.2, we can rewrite the right-handside of (B.1) as
@@
$\qquad=\left\langle\nabla\log\mathbb{P}(z\in\mathscr{Z},y\in\mathscr{Y},o\in\mathscr{O}{\,|\,}x,\theta_{t}),f_{\theta_{t+1}}-f_{\theta_{t}}\right\rangle-\mathrm{KL}\big{(}\mathbb{P}(\cdot,\cdot{\,|\,}x,\theta_{t})\|\mathbb{P}(\cdot,\cdot{\,|\,}x,\theta_{t+1})\big{)},$ (B.16)
where the third line is implied by Lemma B.1, and the last equality follows from (B.1). Plugging (B.1) into (B.1), we have
@@
$\qquad-\mathrm{KL}\big{(}\mathbb{P}(\cdot,\cdot{\,|\,}x,\theta_{t})\|\mathbb{P}(\cdot,\cdot{\,|\,}x,\theta_{t+1})\big{)}.$ (B.17)
Combining (B.1), (B.11), and (B.1), we have
@@
$\qquad-\mathrm{KL}\big{(}\mathbb{P}(\cdot,\cdot{\,|\,}x,\theta_{t})\|\mathbb{P}(\cdot,\cdot{\,|\,}x,\theta_{t+1})\big{)}.$ (B.18)
Putting (B.1) and (B.1) together, we obtain
@@
B.2 Additional Theoretical Results
Theorem B.3.

### v2

@@ 3.2 Bootstrapping Reinforced Thinking Process
This lemma is equivalent to the non-negativity of the KL divergence. For a detailed proof, please refer to Appendix B.1. ∎
@@
Suppose Assumption 3.2 holds. Given that $\mathcal{L}$ is concave and $\theta^{*}=\mathop{\mathrm{argmax}}_{\theta}\mathcal{L}(\theta)$, we have
@@
See Appendix C.1 for a detailed proof. ∎
Theorem 3.3 establishes a convergence rate of $1/T$, where $T$ represents the number of iteration steps. In our analysis, we link our algorithm to classical mirror descent algorithms (Nemirovskij and Yudin, 1983; Bubeck et al., 2015), which have also been utilized in recent studies of policy optimization algorithms (Agarwal et al., 2021; Cai et al., 2020; Zhong and Zhang, 2024). Additionally, our results rely on a concave assumption. This assumption is crucial for proving global convergence results, and we can still obtain a weaker guarantee for stationary points even without this concave condition; see Appendix C.2 for details.
@@ 3.4 Practical Implementation: The Power of Reinforcement Learning
See Appendix B.2 for a detailed proof. ∎
@@ 4.1 Experimental Setups
    BRiTE first obtains $\mathbb{Q}$ in (3.4) through RL, then updates $\theta$ in (3.2) using SFT. The SFT data consists of three components: problems $x$, thinking processes $z_{\mathbb{Q}}(x)$ generated by $\mathbb{Q}$, along with the ground truth answers $y^{*}$.
@@ 4.2 Experimental Result and Analysis
Finally, we extend the evaluation of BRiTE beyond mathematics tasks to assess its performance on code generation. The results, presented in Table 2, demonstrate a consistent trend observed in mathematics tasks: BRiTE outperforms rejection sampling-based algorithms and even surpasses SFT using human-annotated answers. This highlights BRiTE’s versatility and effectiveness across diverse problem domains, showcasing its ability to generalize to a wide range of reasoning tasks. We also remark that rejection sampling-based algorithms require the code datasets to be equipped with correct unit tests, which are unnecessary for BRiTE. For a detailed description of the experimental setup and configuration, refer to Appendix D.2.
@@
4.3 Enhanced Math Reasoning via Expanded Dataset and Advanced Base Model
In the previous subsection, BRiTE’s performance did not demonstrate significant advantages over other algorithms like rejection sampling. This was primarily due to two factors: the instruct models had already undergone post-training, and the dataset size was limited. To overcome these constraints, we have implemented a larger training dataset based on the more advanced Qwen base model.
Specifically, we perform BRiTE on the model Qwen2.5-7B (Team, 2024) with a mixed dataset (the size is around 40K) of RUC-AIBOX/STILL-3-Preview-RL-Data [^2] https://huggingface.co/datasets/RUC-AIBOX/STILL-3-Preview-RL-Data, MATH, and historical AIME problems (excluding AIME 2024). We evaluated models trained using the RS baseline, the $\psi$-update and $\theta$-update of BRiTE on a diverse set of reasoning benchmarks, including six challenging math and science reasoning benchmarks: Math-500, Minerva Math, Olympiad Bench, AIME24, AMC23, and GPQA Diamond. Here MATH500 contains 500 competition-level math problems drawn from the MATH dataset (Hendrycks et al., 2021), spanning subjects from algebra to precalculus and accompanied by solutions . Minerva Math comprises 272 undergraduate-level quantitative problems in physics, chemistry, biology, economics, and mathematics, designed to test advanced multi-step reasoning (Lewkowycz et al., 2022). OlympiadBench (He et al., 2024) is a collection of 8,952 Olympiad-caliber problems in math and physics (with 57% including diagrams) provided in both English and Chinese, each with a detailed solution . Drawing from its extensive entries, we selected 674 open‑ended, competition‑style, text‑only problems from OlympiadBench. AMC23 (40 problems) and AIME24 (30 problems) are benchmarks based on recent AMC 12 (2023) and AIME (2024) contests (Mathematical Association of America, 2023, 2024), representing high school-level competition questions of moderate and high difficulty, respectively . Finally, GPQA Diamond (Rein et al., 2024) consists of 198 graduate-level scientific questions verified by experts. Each of these datasets challenges models with a different profile of difficulty and domain coverage, collectively evaluating a broad spectrum of mathematical reasoning capabilities. We evaluate the pass@1 accuracy (64 sampling times for AIME24 and AMC23 and 8 sampling times for the others) for each benchmark.
Method MATH500 Minerva Math OlympiadBench AIME24 AMC23 GPQA Diamond
44.1 12.9 16.1 0.9 10.1 25.9
RS 54.3 21.0 23.1 5.6 31.6 26.9
BRiTE ($\psi$-update) 79.1 35.0 35.7 14.3 57.7 28.5
BRiTE ($\theta$-update) 76.9 40.6 37.0 14.4 57.1 29.8
BRiTE-iter-2 ($\psi$-update) 80.6 41.3 37.3 14.3 57.9 29.9
BRiTE-iter-2 ($\theta$-update) 78.2 39.8 37.9 15.3 56.4 30.1
Table 2: Performance comparison across different reasoning benchmarks.
Figure 4: Left: Training dynamics of BRiTE with an external verifier and rejection sampling. Right: Mean accuracy of benchmark scores of models trained by BRiTE with an external verifier and reject sampling during the training process.
As shown in Table 2, BRiTE with an external verifier can improve reject sampling (RS) significantly. Specifically, on MATH500, Minerva Math and AMC23, BRiTE improve upon RS by over 15 points in accuracy, showcasing the strong impact of RL-bootstrapped rationales. Moreover, we also show that the iterative training version of BRiTE leads to some improvements, though the training in iteration 1 reaches a plateau that may limit further gains. Figure 4 (left) shows that BRiTE exhibits faster and more stable training loss reduction compared to rejection sampling, indicating more effective use of feedback during optimization. Figure 4 (right) further confirms this trend across benchmarks: BRiTE consistently achieves higher mean accuracy throughout training. These results indicate that BRiTE not only improves final performance, but also accelerates the acquisition of reasoning capabilities during training.
@@ Acknowledgments
Appendix A Author Contributions
This work stems from the valuable contributions and close collaboration of all authors. The first seven authors and ZW are the core contributors to this project, all participating in discussions of problem formulation and theory, as well as experimental implementation. In particular, HZ and ZW lead the project, primarily propose the methodology, derive the theoretical results, guide experimental progress, contribute to early baselines, and write the paper. YY, SZ, and XX mainly contribute to the initial implementation of BRiTE without verifier reward for Llama, Gemma, and Mistral models, while YL, YZ, and ZL mainly contribute to BRiTE with verifier reward for larger datasets and Qwen base models. In addition, YL, YZ, and ZL make key contributions to direct preference learning, rejection sampling, and code generation, respectively. Other authors also make significant contributions to this work, providing computational resources and offering suggestions for theoretical analysis, experiment design, and paper writing.
Appendix B Missing Proofs in the Main Paper
B.1 Proof of Lemma 3.1
@@ Appendix B Missing Proofs in the Main Paper
$\mathbb{E}_{w\sim q(\cdot)}[\log P_{w}-\log q(w)]-\log\sum_{w\in\mathcal{W}}P_{w}$ $=\mathbb{E}_{w\sim q(\cdot)}\Big{[}\log\frac{P_{w}}{\sum_{w^{\prime}\in\mathcal{W}}P_{w^{\prime}}}-\log q(w)\Big{]}=-\mathrm{KL}(q\|\widetilde{p})\leq 0,$
@@
Appendix C Convergence Results
C.1 Proof of Theorem 3.3
@@ Appendix C Convergence Results
$=A(x,\theta^{\prime})-A(x,\theta)+\left\langle\mathbb{E}_{(z,y)\sim\mathbb{P}(\cdot,\cdot{\,|\,}x,\theta)}[\mathcal{K}((x,z,y),\cdot)],f_{\theta}-f_{\theta^{\prime}}\right\rangle,$ (C.1)
@@
$=\sum_{(z,y)\in\mathcal{Z}\times\mathcal{Y}}\mathbb{P}(z,y{\,|\,}x,\theta)\cdot\mathcal{K}((x,z,y),\cdot),$ (C.2)
where $\nabla$ is a functional gradient. Combining (C.1) and (C.1), we have
@@
which finishes the proof of Lemma C.1. ∎
Lemma C.2.
@@
$=\frac{\sum_{(z,y,o)\in\mathscr{Z}\times\mathscr{Y}\times\mathscr{O}}\big{(}\mathbb{P}(o{\,|\,}x,z,y)\cdot\nabla\mathbb{P}(z,y{\,|\,}x,\theta)\big{)}}{\sum_{(z,y,o)\in\mathscr{Z}\times\mathscr{Y}\times\mathscr{O}}\mathbb{P}(z,y,o{\,|\,}x,\theta)},$ (C.3)
@@
$=\mathbb{P}(z,y{\,|\,}x,\theta)\cdot[\mathcal{K}((x,z,y),\cdot)-\nabla A(x,\theta)].$ (C.4)
Combining (C.1) and (C.1), we have
@@
$=\mathbb{E}_{(z,y)\sim\mathbb{Q}(\cdot,\cdot{\,|\,}x,\theta)}[\mathcal{K}((x,z,y),\cdot)]-\nabla A(x,\theta).$ (C.5)
This finishes the proof of Lemma C.2. ∎
@@
$=\mathop{\mathrm{argmax}}_{\theta}\Big{\{}\left\langle\mathbb{E}_{(z,y)\sim\mathbb{Q}(\cdot,\cdot{\,|\,}x,\theta_{t})}[\mathcal{K}((x,z,y),\cdot)],f_{\theta}-f_{\theta_{t}}\right\rangle+A(x,\theta_{t})-A(x,\theta)\Big{\}},$ (C.6)
@@
Combining (C.1) and Lemma C.2, we have
$\theta_{t+1}$ $=\mathop{\mathrm{argmax}}_{\theta}\Big{\{}\left\langle\nabla\log\mathbb{P}(z\in\mathscr{Z},y\in\mathscr{Y},o\in\mathscr{O}{\,|\,}x,\theta_{t})+\nabla A(x,\theta_{t}),f_{\theta}-f_{\theta_{t}}\right\rangle+A(x,\theta_{t})-A(x,\theta)\Big{\}}.$ (C.7)
@@
$\nabla A(\theta_{t+1})-\nabla A(\theta_{t})=\nabla\log\mathbb{P}(z\in\mathscr{Z},y\in\mathscr{Y},o\in\mathscr{O}{\,|\,}x,\theta_{t}).$ (C.8)
@@
where the last equality follows from Lemma C.1. This is equivalent to
@@
$\qquad=\underbrace{\langle\nabla\log\mathbb{P}(z\in\mathscr{Z},y\in\mathscr{Y},o\in\mathscr{O}{\,|\,}x,\theta_{t}),f_{\theta^{*}}-f_{\theta_{t+1}}\rangle}_{(\star)}+\mathrm{KL}\big{(}\mathbb{P}(\cdot,\cdot{\,|\,}x,\theta_{t})\|\mathbb{P}(\cdot,\cdot{\,|\,}x,\theta_{t+1})\big{)},$ (C.9)
where the last equality follows from (C.8). For the Term ()($\star$)( ⋆ ), we have
@@
$=\underbrace{\left\langle\nabla\log\mathbb{P}(z\in\mathscr{Z},y\in\mathscr{Y},o\in\mathscr{O}{\,|\,}x,\theta_{t}),f_{\theta^{*}}-f_{\theta_{t}}\right\rangle}_{\mathrm{(I)}}+\underbrace{\left\langle\nabla\log\mathbb{P}(z\in\mathscr{Z},y\in\mathscr{Y},o\in\mathscr{O}{\,|\,}x,\theta_{t}),f_{\theta_{t}}-f_{\theta_{t+1}}\right\rangle}_{\mathrm{(II)}}.$ (C.10)
@@
$\log\mathbb{P}(z\in\mathscr{Z},y\in\mathscr{Y},o\in\mathscr{O}{\,|\,}x,\theta_{t+1})$ (C.12)
@@
$\mathbb{Q}(z,y,o{\,|\,}x,\theta)=\frac{\mathbb{P}(z,y,o{\,|\,}x,\theta)}{\sum_{(z^{\prime},y^{\prime},o^{\prime})\in\mathscr{Z}\times\mathscr{Y}\times\mathscr{O}}\mathbb{P}(z^{\prime},y^{\prime},o^{\prime}{\,|\,}x,\theta)}=\frac{\mathbb{P}(z,y,o{\,|\,}x,\theta)}{\mathbb{P}(z\in\mathscr{Z},y\in\mathscr{Y},o\in\mathscr{O}{\,|\,}x,\theta)}.$ (C.13)
@@
$=\sum_{(z,y,o)\in\mathscr{Z}\times\mathscr{Y}\times\mathscr{O}}\mathbb{Q}(z,y,o{\,|\,}x,\theta_{t})\cdot\log\frac{\mathbb{P}(z,y,o{\,|\,}x,\theta_{t})}{\mathbb{Q}(z,y,o{\,|\,}x,\theta_{t})},$ (C.14)
where the last equality is implied by the definition of $\mathbb{Q}(z,y,o{\,|\,}x,\theta)$ in (C.13). Combining (LABEL:eq:l:t+1) and (C.1), we have
@@
$\qquad=\sum_{(z,y,o)\in\mathscr{Z}\times\mathscr{Y}\times\mathscr{O}}\mathbb{Q}(z,y,o{\,|\,}x,\theta_{t})\cdot\log\frac{\mathbb{P}(z,y{\,|\,}x,\theta_{t+1})}{\mathbb{P}(z,y{\,|\,}x,\theta_{t})},$ (C.15)
where the third equality follows from the non-negativity of KL-divergence, and the last equality is implied by (3.1). By Assumption 3.2, we can rewrite the right-handside of (C.1) as
@@
$\qquad=\left\langle\nabla\log\mathbb{P}(z\in\mathscr{Z},y\in\mathscr{Y},o\in\mathscr{O}{\,|\,}x,\theta_{t}),f_{\theta_{t+1}}-f_{\theta_{t}}\right\rangle-\mathrm{KL}\big{(}\mathbb{P}(\cdot,\cdot{\,|\,}x,\theta_{t})\|\mathbb{P}(\cdot,\cdot{\,|\,}x,\theta_{t+1})\big{)},$ (C.16)
where the third line is implied by Lemma C.1, and the last equality follows from (C.1). Plugging (C.1) into (C.1), we have
@@
$\qquad-\mathrm{KL}\big{(}\mathbb{P}(\cdot,\cdot{\,|\,}x,\theta_{t})\|\mathbb{P}(\cdot,\cdot{\,|\,}x,\theta_{t+1})\big{)}.$ (C.17)
Combining (C.1), (C.11), and (C.1), we have
@@
$\qquad-\mathrm{KL}\big{(}\mathbb{P}(\cdot,\cdot{\,|\,}x,\theta_{t})\|\mathbb{P}(\cdot,\cdot{\,|\,}x,\theta_{t+1})\big{)}.$ (C.18)
Putting (C.1) and (C.1) together, we obtain
@@
C.2 Additional Theoretical Results
Theorem C.3.